A 3.5

3.5 g sample of sodium hydrogen carbonate is added to a solution of acetic acid weighing 10.5

10.5 g. The two substances react, releasing carbon dioxide gas to the atmosphere. After the reaction, the contents of the reaction vessel weigh 12.4

12.4 g. What is the mass of carbon dioxide released during the reaction?


1
Expert's answer
2021-10-04T05:11:18-0400

m(NaHCO3) = 3.5 g

m(CH3COOH) = 10.5 g

m(reaction vessel) = m(CH3COONa + H2O) = 12.4 g

m(CO2) = unknown


Solution:

Balanced chemical equation:

NaHCO3(aq) + CH3COOH(aq) ⟶ CH3COONa(aq) + H2O(l) + CO2(g)


According to the law of conservation of mass, the mass of the products in a chemical reaction must equal the mass of the reactants.

Hence,

m(NaHCO3) + m(CH3COOH) = m(CH3COONa + H2O) + m(CO2)

m(CO2) = 3.5 g + 10.5 g - 12.4 g = 1.6 g

m(CO2) = 1.6 g


Answer: 1.6 grams of carbon dioxide (CO2) is released during the reaction.